OpenStudy (anonymous):

How do I solve (sqrt)2x+3 = x ?

hartnn (hartnn):

its \(\sqrt{2x+3}=x\) right ? square both sides ,what u get ?

OpenStudy (anonymous):

2x +3 = x^2 ?

OpenStudy (anonymous):

Then I solve from there?

hartnn (hartnn):

yes, you know how to solve that quadratic equation ?

OpenStudy (anonymous):

I'm not sure..

OpenStudy (anonymous):

It looks easy! I don't know what I'm missing..

hartnn (hartnn):

we have x^2-2x-3 =0 ok ? can you tell me 2 numbers whose sum is -2 and product is -3 ??

hartnn (hartnn):

hint : one number is positive and other is negative

OpenStudy (anonymous):

-3 times 1?

hartnn (hartnn):

correct! so we split the -2x into -3x+1x we have now x^2-2x+x-3 =0 know how to proceed further ?

OpenStudy (anonymous):

combine like terms?

hartnn (hartnn):

factor out 'x' from the 1st 2 terms x^2-2x what u get ?

OpenStudy (anonymous):

wait, I don't understand. What are we doing with the x's?

hartnn (hartnn):

we are trying to factor x^2-2x-3=0 and get and answer like (x-a)(x+b) = 0

OpenStudy (anonymous):

would it be 0 = (x-3)(x+1) ?

hartnn (hartnn):

correct! good :)

hartnn (hartnn):

(x-3)(x+1) = 0 so x-3=0 , x = .... ? x+1 =0, x= ... ?

hartnn (hartnn):

correct! good :)

OpenStudy (anonymous):

x=3 and x=-1

OpenStudy (anonymous):

Thank you so much! I should have known this. I appreciate it. (:

hartnn (hartnn):

you are most welcome ^_^